How to find resultant vector?
Such that a person walk 10 km east and 20 km north what is the resultant of this?

Grade:12th pass

1 Answers

Shubham pandey
6 years ago
R=√a^2+b^2+2abcos90. Because they are perpendicular to each other.So=√10^2+20^2+0=√500=10√5 is the resultant